# On the interpretation of quantum theory as games between physicists and nature played in Minkowski spacetime

### v1

@@ 2.2 Standardization of Bob’s actions in alternating spacetime games
Without loss of generality, we can put constraints on Bob’s actions in an alternating game to facilitate their interpretation as contexts, like so:
  • Each edge from a node played by Bob to a node played by Alfred is labeled with a set (the context) of information sets (the settings) played by Alfred.
@@
In other words, Bob’s actions are contexts, and for each context, there are edges to each measurement setting in that context, labeled with that context.
It is always possible to rename Bob’s actions such that this is the case: Given a node $N$ played by Bob and an action $c\in\chi(N)$, we rename $c$ to $\{\iota(M)|M\in\mathcal{N}\wedge\sigma(N,M)=c\}$. Because of alternating condition 6, the renaming is bijective.
We show below a correspondence table between the two frameworks.
@@
An information set played by Bob a set of events, which is the left-hand-
side of an enabling relation
@@
An edge from Bob to Alfred Bob picks a measurement context and
carries out a measurement in this context.
@@
An edge from Alfred to Bob An outcome $o$ is obtained for a specific
measurement $x$, and is part of enabling
conditions $t$ for another measurement.
The label is $t(x)=o$
@@ 2.4 Natural cover
Each spacetime game (not necessarily fulfilling the alternating property) has a natural cover that corresponds to all combinations of measurements that are performed in one of the complete histories:
@@
Given a spacetime game $\mathcal{G}$ with nature player Alfred, the natural cover $\mathcal{C}$ is defined as follows: For each outcome $z\in Z$, we take the subset $C_{z}=\{i\in\textsf{domain}(z)|\rho(i)=\text{Alfred}\}$ of its domain of definition with only and exactly the information sets that are played by Alfred. Then the natural cover is $\mathcal{C}=\{C_{z}|z\in Z\}$. Note that duplicates are implicitly eliminated because of the definition of a set.
For example, the natural cover of the game shown in Figure 1 and Figure 2 is $\{\{XW\},\{XZ\},\{YW\},\{YZ\}\}$.
The natural cover of the spacetime game shown in Figure 3 and Figure 4 is $\{\{XY\},\{XZ\},\{YZ\}\}$.
@@
This natural cover will be used to define the corresponding measurement scenario for any alternating game.
@@ 2.5 The mapping
  • Bob’s actions (which label edges from Bob to Alfred) are contexts. Alfred’s nodes within the same information set are measurements with the same measurement setting but in different contexts;
  • Alfred’s actions (which label edges to Bob) are labeled by measurement outcomes.
Definition 8 (Causal contextuality scenario associated with a spacetime game).
Given an alternating spacetime game $\mathcal{G}$ with nature player $\text{Alfred}\in\mathcal{P}$, we define the equivalent causal measurement scenario $\Sigma(\mathcal{G})=(X,O,\vdash)$ as follows:
@@
  • for each $x\in X$, we define that $t\vdash x$ where
    $t=\{(\iota(M),\sigma(M,\textsf{parent}(x)))|M\smile\textsf{parent}(x)\}$
    Note that when $\textsf{parent}(x)=\{\}$, this amounts to {}$\{\}\vdash x$. Also note that it means that $\sigma(M,\textsf{parent}(x))=t(y)$ for any $M\in y$.
  • the cover $\mathcal{C}$ is the natural cover of $\mathcal{G}$.
We use the abbreviation $\textsf{enabled}(t)=\{x|t\vdash x\}$, in particular for restricting the cover $\mathcal{C}$ to $\mathcal{C}\cap\textsf{enabled}(t)$ which we define as
$\mathcal{C}\cap\textsf{enabled}(t)=\{C\cap\textsf{enabled}(t)|C\in\mathcal{C}\}\setminus\{\{\}\}$
The property that alternating spacetime games have unique causal bridges (Dupuy’s terminology for enabling relations) translates to the same for the associated causal contextuality scenarios.
Definition 9 (Unique causal bridges).
Let $(X,O,\vdash,\mathcal{C})$ be a causal contextuality scenario. We say that the causal contextuality scenario has unique causal bridges if every measurement is on the right-hand-side of exactly one enabling relation (possibly with an empty left-hand side).
@@
The causal contextuality scenario associated with an alternating spacetime game always has unique causal bridges.
@@
This is true by construction, since only one enabling relation is defined for every $x\in X$. ∎
Note that it is always possible to turn a causal contextuality scenario into a causal contextuality scenario that has unique causal bridges by duplicating and renaming measurements to make the right-hand sides of $\vdash$ unique, although it is not neutral in terms of semantics. This amounts to saying that, when a measurement experiment is performed (also with multiple contexts possible), the set of events that enabled this measurement is known to the experimenter, which is reasonable to assume in a local lab. However, a discussion of extending the isomorphism to non-unique causal bridges is given in Section 3. We focus our attention specifically to unique causal bridges in this paper.
@@
Given a canonical causal contextuality scenario $(X,O,\vdash,\mathcal{C})$, we define the associated spacetime game with imperfect information as follows:
@@
  • Nodes played by Bob: $\mathcal{N}_{B}=\{t|\exists x,t\vdash x\}$
  • Nodes played by Alfred: $\mathcal{N}_{A}=\{(x,c)|x\in X,t\vdash x,c\in\mathcal{C}\cap\textsf{enabled}(t)\}$
@@
  • $\forall(x,c)\in\mathcal{N}_{A},\rho((x,c))=\text{Alfred}$
@@
  • Actions by Bob: $\mathcal{A}_{B}=\{\mathcal{C}\cap\textsf{enabled}(t)|t\in\mathcal{N}_{B}\}$
  • Actions by Alfred: $\mathcal{A}_{A}=\bigcup_{x}O_{x}$
@@
  • $\forall(x,c)\in\mathcal{N}_{A},\chi(x,c)=O_{x}$
  • $\forall t\in\mathcal{N}_{B},\chi(t)=\mathcal{C}\cap\textsf{enabled}(t)$
  • $\mathcal{R}_{A}=\{(t,(x,c))|t\vdash x\}$
  • $\mathcal{R}_{B}=\{((x,c),t)|x\in\textsf{domain}(t)\}$
@@
  • $\forall(t,(x,c))\in\mathcal{R}_{A},\sigma(t,(x,c))=c$
  • $\forall((x,c),t)\in\mathcal{R}_{B},\sigma((x,c),t))=t(x)$
  • Alfred’s information sets: $\mathcal{I}_{A}=\{(x,c)|x\in X\}\equiv X$
@@
The spacetime game associated with a canonical causal contextuality scenario with unique causal bridges is always alternating.
@@
The first property is true by construction.
The second property, that the graph is bipartite, holds by construction, as edges always connect a node in $\mathcal{N}_{A}$ to a node in $\mathcal{N}_{B}$ or vice versa.
The third property, that Bob’s information sets are singletons, also holds by construction since $\mathcal{I}_{B}=\mathcal{N}_{B}$.
The fourth property is that the nodes in the same information set of Alfred have a unique parent. These nodes are necessarily of the form $(x,c_{1})$, $(x,c_{2})$ for some measurement setting $x$ and contexts $c_{1}$ and $c_{2}$. By construction, the parent node of both of them must be some $t$ such that $t\vdash x$, and thus must be the same because we assume that the causal contextuality scenario has unique causal bridges.
The fifth property is that Bob’s information sets always have one successor. This is true because every one of them corresponds to the left hand side of one or more enabling relations, and the right-hand side of any of them gives a successor.
The sixth property is that all outgoing edges of the nodes $((x,c))_{c}$ in one of Alfred’s information sets $x$ have the same outgoing edges. This is true from the definition $\mathcal{R}_{B}=\{((x,c),t)|x\in\textsf{domain}(t)\}$ as well as of $\forall((x,c),t)\in\mathcal{R}_{B},\sigma((x,c),t))=t(x)$. Indeed, there is no dependency of any edge and label on the context.
The seventh property is that two distinct nodes played by Bob cannot have the same causal bridges. If we take two of these nodes, they correspond to two left-hand-sides of enabling relations $t$ and $s$. Let us further assume that the nodes for $t$ and $s$ have the same predecessors. This means by construction that $t$ and $s$ have the same domains, since the predecessors are made of the domain of $t$ plus contexts. Furthermore, if we assume that the labels of the edges from the predecessors are the same for $t$ and $s$, it implies that $t$ and $s$ exactly match on their full domain, also by construction of the labels, which are of the form $t(x)$ and $s(x)$ for each predecessor $x$. Thus, $t=s$ and the fourth property is fulfilled.
For the eigth property, let us take a node played by Bob (the left hand-side $t$ of some enabling relation), and that is connected to two nodes played by Alfred. These must be of the form $(x,c_{1})$ and $(y,c_{2})$ with $t\vdash x$ and $t\vdash y$. Let us also assume that the labels are identical. These labels correspond to the contexts in which $x$ and $y$ are measured, meaning that $x$ and $y$ are performed in the same context $c=c_{1}=c_{2}$. But then, $x$ and $y$ must be different, otherwise the nodes would be the same $(x,c)=(y,c)$.
Thus, the spacetime game associated with a causal contextuality scenario always fulfills the alternating property. ∎
@@
The mapping from alternating spacetime games to canonical causal contextuality scenarios is injective.
@@
We prove that if a spacetime game $\mathcal{G}$ is mapped to a given canonical causal contextuality scenario $(X,O,\vdash,\mathcal{C})$, then $\mathcal{G}$ is necessarily (up to an isomorphism) the game associated with the causal contextuality scenario.
@@
    $\mathcal{P}=\{\text{Alfred},\text{Bob}\}$: this follows from the assumption that the game is alternating (first condition).
@@
    It must be that $\mathcal{I}_{A}=X$ for the original game by construction of the causal contextuality scenario.
@@
    The left hand sides of the enabling relation are, by definition of the mapping from $\mathcal{G}$, exactly those of the form
    $t=\{(\iota(M),\sigma(M,P))|M\smile P\}$
    for each $P$ that is a node played by Bob in $\mathcal{G}$ and that is the parent of $x$ such that $t\vdash x$.
    Because of the alternating property, for any $t$ the corresponding node $P$ must be unique, as two of Bob’s nodes cannot have the same causal bridge (seventh condition). Conversely, any of Bob’s nodes must have a successor information set $x$ (also because of the alternating property, fifth condition) and thus, any of Bob’s nodes in the original game yields the left-hand side of an enabling relation. Thus, the correspondence between Bob’s nodes and the left-hand-sides of the enabling relations is exact: $\mathcal{N}_{B}=\{t|\exists x,t\vdash x\}$.
    This yields the pattern shown on Figure 7. It also follows that $\textsf{enabled}(t)=\textsf{successor}(t)$.
    Figure 7: The graph pattern implied by the definition of $t\vdash x$ in the mapping from games to causal contextuality scenarios. We show the pattern for two events in $t=\{(y,t(y)),(z,t(z))\}$ and two contexts for $x$ for $y$ and for $z$, but this extends to any number of events and any domain size.
@@
    $\mathcal{I}_{B}=\mathcal{N}_{B}$ necessarily because the original game is alternating (third condition).
@@
    Whenever $t\vdash x$ with $t$ being one of Bob’s nodes and $x$ one of Alfred’s information sets, the parent of all the nodes in $x$ must be $t$, because it exists and is unique over the entire information set according to the alternating property (fourth condition). It follows that edges from Bob to Alfred are of the form $(t,(x,c))$. Thus, $\mathcal{R}_{A}=\{(t,(x,c))|t\vdash x\}$ where the allowed values for $c$, which index the nodes in information set $x$, $((x,c))_{c}$, must still be defined.
@@
    $\forall(t,(x,c))\in\mathcal{R}_{A},\sigma(t,(x,c))=c$ follows from the renaming convention of the labels from Bob to Alfred. According to this convention, for any node $t$ played by Bob, there are as many outgoing edges from $t$ as there are pairs $(x,d)$ such that $x\in d$, $d\in\chi(t)$, and $\sigma(t,(x,d))=d$. Thus $d$ also indexes the nodes in the information set $x$ and can be identified with $c$, and the edge going from $t$ to node $(x,c)$ must be labeled with $\sigma(t,(x,c))=c$.
@@
    We now show that $\chi(t)=\mathcal{C}\cap\textsf{enabled}(t)$. This amounts to showing that $\chi(t)=\mathcal{C}\cap\textsf{successor}(t)$
    Because of the naming convention of Bob’s action, we know that the labels in $\chi(t)$ are sets of information sets that are successors of $t$.
    Recall that $\mathcal{C}$ is by construction of the causal contextuality scenario a natural cover. Let us take an element $c$ in $\mathcal{C}\cap\textsf{successor}(t)$. By definition, $c$ is not the empty set. There is a corresponding complete history $z$ of the original game such that $\forall i\in\textsf{successor}(t),i\in c\iff z(i)\neq\bot$. This means that $c$ contains exactly the information sets that are activated by $t$ in $z$, and it there is at least one.
@@
    The converse is also true. If we take an action $c\in\chi(t)$, it labels the edges to all the nodes $(x,c)$ such that $x\in c$ because of the naming convention, and it is not empty because at least one successor exists. Thus, there exists a complete history assigning $c$ to $t$ (because the game is clean as has no unused nodes) and assigning each $x\in c$ to some value, and no value for the other successors of $t$. The active information sets of Alfred in this complete histories form a set that is an element $\hat{c}$ of $\mathcal{C}$, the natural cover, which is a superset of $c$. Thus $c=\hat{c}\cap\textsf{successor}(t)$, and $c$ is not the empty set. Thus, $c\in\mathcal{C}\cap\textsf{successor}(t)$.
    We have thus proven that $\chi(t)=\mathcal{C}\cap\textsf{enabled}(t)$.
    We also note that there cannot be any duplicates, i.e., there cannot be distinct nodes in $x$ indexed by the same $c\in\chi(t)$, because of the eight condition of the alternating property.
@@
    We have shown that the nodes in an information set $x$ are indexed by $\chi(t)$ where $t\vdash x$, so that $\mathcal{N}_{A}=\{(x,c)|x\in X,t\vdash x,c\in\chi(t)\}$. Also, we have shown that $\chi(t)=\mathcal{C}\cap\textsf{enabled}(t)$. Thus $\mathcal{N}_{A}=\{(x,c)|x\in X,t\vdash x,c\in\mathcal{C}\cap\textsf{enabled}(t)\}$.
@@
    Since we have established the pattern on Figure 7, it follows that edges from Alfred to Bob connect each information set in the domain of $t$ to $t$. Thus, the edges from Alfred to Bob are given by $\mathcal{R}_{B}=\{((x,c),t)|x\in\textsf{domain}(t)\}$. All these edges must be present because of the alternating property’s (sixth condition) implying that the nodes in an information must have the same successors.

### v2

@@ 7. BA2
  8. AB1
    $\forall i\in\rho^{-1}(\text{Alfred}),\forall N,M\in i,\textsf{successors}(N)=\textsf{successors}(M)$
    $\forall i\in\rho^{-1}(\text{Alfred}),\forall N,M\in i,\forall T\in\textsf{successors}(i),\sigma(M,T)=\sigma(N,T)$
  9. AB2
    Two distinct nodes played by Bob cannot have the same causal bridge.
    $\forall N,M\in\rho^{-1}(\text{Bob}),$
    $(\textsf{predecessors}(N)=\textsf{predecessors}(M)\wedge$
    $\forall P\in\textsf{predecessors}(N),\sigma(P,N)=\sigma(P,M))\implies N=M$
Remark 1.
The height of an alternating spacetime game is always even, nodes at even depths (0, 2, …) are played by Bob, and nodes at odd depths (1, 3, …) are played by Alfred.
3.2 Non-locality as a special case of contextuality
AB2 amounts to saying that we assume that the measurements carried out by Bob under the same enabling circumstances are grouped into a unique experiment where Bob has multiple contexts to choose from. Note that this does not cause any loss of generality because this amounts to saying that non-locality is a special case of contextuality. See for example Figure 7 (which satisfies AB2), which groups the local experiments of Figure 1 (which does not satisfy AB2) into a single one.
The purpose of alternating spacetime games is to characterize a subset of spacetime games that is isomorphic to causal contextuality scenarios with good covers. When using the spacetime game framework for discussing non-locality, it is more desirable to consider the (factored) game in Figure 1 rather than the (expanded) game in Figure 7, knowing that the factored game can be considered equivalent to the same causality contextuality scenario as the one associated with the expanded game by the mapping.
@@ 3.3 Standardization of Bob’s actions in alternating spacetime games
Without loss of generality, we can put constraints on Bob’s actions in an alternating spacetime game to facilitate their interpretation as contexts, like so:
  • Each edge from a node played by Bob to a node played by Alfred is labeled with a non-empty set (the context) of information sets (the settings) played by Alfred.
@@
In other words, Bob’s actions are contexts, and for each context, there are edges to each measurement setting in that (possibly empty) context, labeled with that context.
It is always possible to rename Bob’s actions such that this is the case: Given a node $N$ played by Bob and an action $c\in\chi(N)$, we rename $c$ to $\{\iota(M)|M\in\mathcal{N}\wedge\sigma(N,M)=c\}$. Because of BA2, the renaming is bijective.
We show below a correspondence table between the two frameworks in Table 1.
@@
An information set played by Bob A set of events, which is the left-hand-side of an enabling relation
@@
An edge from Bob to Alfred Bob picks a measurement context and carries out a measurement in this context.
@@
An edge from Alfred to Bob An outcome $o$ is obtained for a specific measurement $x$, and is part of enabling
conditions $t$ for another measurement. The label is $t(x)=o$
A complete history A complete history
A history with an even number of defined decisions A history
Table 1: The correspondence of terminology between alternating spacetime games and causal contextuality scenarios.
@@ 3.4 Causal contextuality scenarios
In this paper, we consider “clean” scenarios in the sense that there are no unused measurement settings, i.e., settings that cannot be enabled under any circumstances under the considered cover. If a causal contextuality scenario contains an unused setting, then it can be removed without affecting its semantics. This is analogous to our assumption that spacetime games do not have unused information sets.
Furthermore, we restrict our attention in this paper to covers that satisfy the antichain property. It is, however, possible to extend the mapping naturally even if the antichain property does not hold (e.g., by allowing unused actions at Bob’s nodes).
@@ 3.5 Natural cover
Each spacetime game (not necessarily alternating) has a natural cover that corresponds to all combinations of measurements that are performed in one of the complete histories:
@@
Given a spacetime game $\mathcal{G}$ with nature player Alfred, the natural cover $\mathcal{C}$ is defined as follows: For each outcome $z\in\mathcal{Z}$, we take the subset $C_{z}=\{i\in\textsf{domain}(z)|\rho(i)=\text{Alfred}\}$ of its domain of definition with only and exactly the information sets that are played by Alfred. Then the natural cover is $\mathcal{C}=\{C_{z}|z\in\mathcal{Z}\}$. Note that duplicates are implicitly eliminated because of the definition of a set.
For example, the natural cover of the game shown in Figure 1 (or 7) and Figure 2 (or 8) is $\{\{XW\},\{XZ\},\{YW\},\{YZ\}\}$, which is the cycle of rank 4.
The natural cover of the spacetime game shown in Figure 3 and Figure 4 is $\{\{XY\},\{XZ\},\{YZ\}\}$, which is the cycle of rank 3.
@@
This natural cover will be used to define the corresponding measurement scenario for any alternating spacetime game.
@@ 3.6 The mapping
3.6.1 Causal contextuality scenario associated with an alternating spacetime game
@@
  • Bob’s actions (which label edges from Bob to Alfred) are contexts. Alfred’s nodes within the same information set are measurements with the same measurement setting but in different contexts and possibly in different causal pasts (i.e., different parent nodes);
  • Alfred’s actions (which label edges to Bob) are the measurement outcomes.
We now give a formal definition of the above:
Definition 7 (Causal contextuality scenario associated with an alternating spacetime game).
Given an alternating spacetime game $\mathcal{G}$ with nature player $\text{Alfred}\in\mathcal{P}$, we define the equivalent causal contextuality scenario $\Sigma(\mathcal{G})=(X,O,\vdash,\mathcal{C})$ as follows:
@@
  • for each $x\in X$, and for each $P\in\textsf{predecessors}(x)$, we define that $t_{P}\vdash x$ where
    $t_{P}=\{(\iota(M),\sigma(M,P))|M\smile P\}$
  • additionally, for each $x\in X$, if there is a node in $x$ with no parent, we also define {}$\{\}\vdash x$.
  • $\mathcal{C}$ is the natural cover of $\mathcal{G}$.
We must prove that the enabling relations have consistent left-hand sides for it to be a valid causal contextuality scenario.
@@
The causal contextuality scenario associated with an alternating spacetime game always has consistent sets of events on the left-hand sides of enabling relations.
@@
AB1 implies that two nodes in $y$, which is played by Alfred, have the same labels and destination nodes. As a consequence, if we have two nodes $M,N\in y$, it follows that $\sigma(M,P)=\sigma(N,P)$, leading to only one assignment for $t(y)$. ∎
Remark 2.
It follows from the lemma, given some $x\in X$ and one of its predecessors $P$, that $\sigma(M,P)=t_{P}(y)$ for any $M\in y$. See also Figure 9 for a visual of this pattern.
Figure 9: The graph pattern implied by the definition of $t\vdash x$ in the mapping from games to causal contextuality scenarios. We show the pattern for two events in $t=\{(y,t(y)),(z,t(z))\}$ two contexts for $y$ and $z$, and four contexts (among which two under causal bridge $t$) for $x$, but this extends to any number of events and any domain size.
Lemma 2.
$P\mapsto t_{P}$ is a bijection between the set of nodes played by Bob and the left-hand sides of the mapped enabling relations. More precisely, for any $x\in X$, the left-hand sides of the enabling relations to $x$ in the causal contextuality scenario are in bijection with the predecessors of $x$ in the alternating spacetime game.
Proof.
Every $P$ leads to at least one enabling relation with left-hand-side $t_{P}$ because it has at least one child (property EVEN). It follows from the property AB2 that $t$ is injective. Since the codomain is defined as the range, the mapping is also surjective. ∎
We use the abbreviation $\textsf{enabled}(t)=\{x\in X|t\vdash x\}$ and use this notation in particular for restricting the cover $\mathcal{C}$ to the “local cover at t” denoted $\mathcal{C}_{t}$ which we define as
$\mathcal{C}_{t}=\{C\cap\textsf{enabled}(t)|C\in\mathcal{C}\}\setminus\{\emptyset\}$
Note that $\mathcal{C}_{t}$ cannot be empty because there are no unused settings, and it also does not contain the empty set.
3.6.2 Alternating spacetime game associated with a causal contextuality scenario
We now give the mapping in the opposite direction, i.e., how to associate a spacetime game with a causal contextuality scenario.
@@ 3.6.2 Alternating spacetime game associated with a causal contextuality scenario
Given a causal contextuality scenario $(X,O,\vdash,\mathcal{C})$, we define the associated spacetime game with imperfect information obtained like so:
@@
  • Nodes played by Bob: $\mathcal{N}_{B}=\{t|\exists x\in X,t\vdash x\}$
  • Nodes played by Alfred: $\mathcal{N}_{A}=\{(t,x,c)|t\vdash x,c\in\mathcal{C}_{t}\}$
@@
  • $\forall(t,x,c)\in\mathcal{N}_{A},\rho((t,x,c))=\text{Alfred}$
@@
  • Actions by Bob: $\mathcal{A}_{B}=\bigcup_{t\in\mathcal{N}_{B}}\mathcal{C}_{t}$
  • Actions by Alfred: $\mathcal{A}_{A}=\bigcup_{x\in X}O_{x}$
@@
  • $\forall(t,x,c)\in\mathcal{N}_{A},\chi(t,x,c)=O_{x}$
  • $\forall t\in\mathcal{N}_{B},\chi(t)=\mathcal{C}_{t}$
  • $\mathcal{R}_{A}=\{(t,(t,x,c))|t\vdash x\}$
  • $\mathcal{R}_{B}=\{((u,x,c),t)|x\in\textsf{domain}(t)\}$
@@
  • $\forall(t,(t,x,c))\in\mathcal{R}_{A},\sigma(t,(t,x,c))=c$
  • $\forall((u,x,c),t)\in\mathcal{R}_{B},\sigma((u,x,c),t))=t(x)$
  • Alfred’s information sets: $\mathcal{I}_{A}=\{(t,x,c)|x\in X\}\equiv X$
@@
  • $\mathcal{Z}$ is as defined in Definition 3 to make the game valid.
@@
The spacetime game associated with a causal contextuality scenario is alternating.
@@
The game is valid by construction. Thus, we need to show that it is alternating.
2-PLAYERS is true by construction since $\rho$ can only take two values.
BIPARTITE holds by construction, as edges always connect a node in $\mathcal{N}_{A}$ to a node in $\mathcal{N}_{B}$ or vice versa.
EVEN holds by construction: any node $(t,x,c)$ played by Alfred has $t$ as its parent node, so all root notes must be played by Bob. Likewise, given a node $t$ played by Bob, there is at least one enabling relation $t\vdash x$ that it was taken from. Furthermore, $\mathcal{C}_{t}$ is never empty, so as a consequence, $t$ has at least one child node.
BOB-S (Bob’s information sets are singletons) also holds by construction since $\mathcal{I}_{B}=\mathcal{N}_{B}$.
BOB-A (no unused action at Bob’s nodes) is true: given a node $t$ played by Bob and an action at this node, i.e., a context $c$ in (non-empty) $\mathcal{C}_{t}$, $c$ is used on the edge from $t$ to $x$ for any $x\in c$.
BA1 (nodes played by Alfred have exactly one parent) is true by the construction of $\mathcal{R}_{A}$ because the parent of $(t,x,c)$ is necessarily $t$.
For BA2, let us take a node played by Bob (the left-hand side $t$ of some enabling relation), and that is connected to two nodes played by Alfred. These must be of the form $(t,x,c_{1})$ and $(t,y,c_{2})$ with $t\vdash x$ and $t\vdash y$. Let us also assume that the labels are identical. These labels correspond to the contexts in which $x$ and $y$ are measured, meaning that $x$ and $y$ are performed in the same context $c=c_{1}=c_{2}$. But then, $x$ and $y$ must be different, otherwise the nodes would be the same $(t,x,c)=(t,y,c)$.
AB1 requires that all outgoing edges of the nodes $((u,x,c))_{u,c}$ in one of Alfred’s information sets $x$ have the same outgoing edges. This is true from the definition $\mathcal{R}_{B}=\{((u,x,c),t)|x\in\textsf{domain}(t)\}$ as well as of $\forall((u,x,c),t)\in\mathcal{R}_{B},\sigma((u,x,c),t))=t(x)$. Indeed, there is no dependency of any edge and label on the context.
AB2 requires that two distinct nodes played by Bob cannot have the same causal bridges. If we take two of these nodes, they correspond to two left-hand sides of enabling relations $t$ and $s$. Let us further assume that the nodes for $t$ and $s$ have the same predecessors. This means by construction that $t$ and $s$ have the same domains since the predecessors are made of the domain of $t$ plus contexts. Furthermore, if we assume that the labels of the edges from the predecessors are the same for $t$ and $s$, it implies that $t$ and $s$ exactly match on their full domain, also by construction of the labels, which are of the form $t(x)$ and $s(x)$ for each predecessor $x$. Thus, $t=s$ and BA1 is fulfilled.
Thus, the spacetime game associated with a causal contextuality scenario is always alternating. ∎
@@
The mapping from alternating spacetime games to causal contextuality scenarios as defined in Definition 7 is injective.
@@
We prove that if a spacetime game $\mathcal{G}$ is mapped to a given causal contextuality scenario $(X,O,\vdash,\mathcal{C})$ according to Definition 7, then $\mathcal{G}$ is necessarily, up to an isomorphism [^8] We mean here up to renaming the player names, the labels, etc without affecting the semantics of the game., the game associated with the causal contextuality scenario according to Definition 8.
@@
    $\mathcal{P}=\{\text{Alfred},\text{Bob}\}$: this follows from property 2-PLAYERS.
@@
    It must be that $\mathcal{I}_{A}=X$ for the original game by construction of the causal contextuality scenario (first bullet point of Definition 7).
@@
    The left-hand sides of the enabling relation are, by definition of the mapping from $\mathcal{G}$, exactly those of the form
    $t(P)=\{(\iota(M),\sigma(M,P))|M\smile P\}$
    for each $P$ that is a node played by Bob in $\mathcal{G}$ and that is the parent of some $M\in x$ such that $t\vdash x$.
    Lemma 2 showed that this is a bijection (because of AB2 and EVEN) and that $\mathcal{N}_{B}$ must therefore match the set of the left-hand sides of the enabling relations: $\mathcal{N}_{B}=\{t|\exists x\in X,t\vdash x\}$.
    As a reminder, this yields the pattern shown in Figure 9. It also follows that $\textsf{enabled}(t)=\textsf{successor}(t)$.
@@
    $\mathcal{I}_{B}=\mathcal{N}_{B}$ necessarily because of BOB-S.
@@
    If we consider an information set $x$, the parent of each node in $x$ must be unique according to BA1. Furthermore, if we consider all enabling relations activating $x$, the parents of nodes in $x$ are exactly and all the values of $t$ such that $t\vdash x$ by definition of the enabling relation in Definition 7. It follows that edges from Bob to Alfred are of the form $(t,(t,x,c))$. Thus, $\mathcal{R}_{A}=\{(t,(t,x,c))|t\vdash x\}$ where the allowed values for $c$, which index the nodes in information set $x$ with parent $t$, $((t,x,c))_{c}$, must still be defined.
@@
    $\forall(t,(t,x,c))\in\mathcal{R}_{A},\sigma(t,(t,x,c))=c$ follows from the renaming convention of the labels from Bob to Alfred. According to this convention, for any node $t$ played by Bob, there are as many outgoing edges from $t$ as there are pairs $(x,d)$ such that $x\in d$, $d\in\chi(t)$, and $\sigma(t,(x,d))=d$. Thus $d$ also indexes the nodes in the information set $x$ and can be identified with $c$, and the edge going from $t$ to node $(t,x,c)$ must be labeled with $\sigma(t,(t,x,c))=c$.
@@
    We now show that $\chi(t)=\mathcal{C}_{t}$. This amounts to showing that $\chi(t)=\{C\cap\textsf{successor}(t)|C\in\mathcal{C}\}\setminus\{\emptyset\}$.
    Because of the naming convention of Bob’s actions, we know that the labels in $\chi(t)$ are sets of information sets that are successors of $t$. Thus $\chi(t)\subseteq\mathcal{P}(\textsf{successor}(t))$,
    Recall that $\mathcal{C}$ is by construction of the causal contextuality scenario a natural cover. Let us take an element $c$ in $\mathcal{C}_{t}$. By definition, $c$ is not the empty set, and $c$ is the intersection of some context $C\in\mathcal{C}$ with the set of successors of $t$. Thus, there is a corresponding complete history $z$ of the original game such that $\forall i\in\textsf{successor}(t),i\in c\iff z(i)\neq\bot$. This means that $c$ contains exactly the information sets that are activated by $t$ in $z$, and because $c$ is not empty, there is at least one.
@@
    The converse is also true. If we take an action $c\in\chi(t)$, it labels the edges to all the nodes $(t,x,c)$ such that $x\in c$ because of the naming convention, and $c$ is not the empty set because of EVEN and BOB-A. Thus, there exists a complete history assigning $c$ to $t$ (otherwise $t$ would be an unused node, contradicting that the game is clean) and assigning each $x\in c$ to some value, and no value for the other successors of $t$. The active information sets of Alfred in this complete history form a set that is an element $C$ of $\mathcal{C}$, the natural cover, which is a superset of $c$. Thus $c=C\cap\textsf{successor}(t)$, and $c$ is not the empty set. Thus, $c\in\mathcal{C}_{t}$.
    We have thus proven that $\chi(t)=\mathcal{C}_{t}$.
    We also note that there cannot be any duplicates, i.e., there cannot be distinct nodes in $x$ with the same parent $t$ and indexed by the same $c\in\chi(t)$, because of BA2.
@@
    We have shown that the nodes in an information set $x$ with parent $t$ (where $t\vdash x$) are indexed by $\chi(t)$, so that $\mathcal{N}_{A}=\{(t,x,c)|x\in X,t\vdash x,c\in\chi(t)\}$. Also, we have shown that $\chi(t)=\mathcal{C}_{t}$. Thus $\mathcal{N}_{A}=\{(t,x,c)|x\in X,t\vdash x,c\in\mathcal{C}_{t}\}$.
@@
    Since we have established the pattern in Figure 9, it follows that edges from Alfred to Bob connect each information set in the domain of $t$ to $t$. Thus, the edges from Alfred to Bob are given by $\mathcal{R}_{B}=\{((u,x,c),t)|x\in\textsf{domain}(t)\}$. All these edges must be present (for all $u$ and $c$ such that $(u,x,c)$ is a node) because of AB1’s implying that the nodes in an information set must have the same successors.